OpenStudy (anonymous):

Which points are 56 units from the point located at 13 on a number line? A. –69 and 69 B. –69 and 43 C. –43 and 69 D.

OpenStudy (anonymous):

wow this is hard....and btw i think is D....lol

OpenStudy (fretje):

is this not more a math question? Natural numbers plus zero and the negative numbers are elements of \[\mathbb{Z}\] The two numbers that have a distance of 56 from 13 are (13+56) = 69 and (13-56) = -43

OpenStudy (anonymous):

-43 and 69